# Multiwavelength coverage of the 2024 periastron passage of PSR B125963 / LS 2883

### v1

@@ Acknowledgments
The authors acknowledge support by the state of Baden-Württemberg through bwHPC. We thank Rajan Chhetri and Jamie Stevens for their assistance during the ATCA observations. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. The Australia Telescope Compact Array (ATCA) is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Gomeroi people as the Traditional Owners of the ATCA observatory site. Some of the observations reported in this paper were obtained with the Southern African Large Telescope (SALT) under program 2024-1-SCI-040 (PI: B. van Soelen).

### v2

@@ Acknowledgments
The authors acknowledge support by the state of Baden-Württemberg through bwHPC. We thank Rajan Chhetri and Jamie Stevens for their assistance during the ATCA observations. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. The Australia Telescope Compact Array (ATCA) is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Gomeroi people as the Traditional Owners of the ATCA observatory site. Some of the observations reported in this paper were obtained with the Southern African Large Telescope (SALT) under program 2024-1-SCI-040 (PI: B. van Soelen). The authors thank the European Space Agency (ESA) for the support in the framework of the PRODEX Programme, PEA 4000120711. The authors wish to acknowledge financial support from the Centre for Astrophysics and Relativity at DCU.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding authors.